Wednesday, April 12, 2000 (DOY 096/19:00:00 to DOY 0103/19:00:00 UTC)
Launch / Days since Launch = Nov 7, 1996 / 1253 days
Start of Mapping / Days since Start of Mapping = April 1, 1999 / 377 days Total Mapping Orbits = 4906
Total Orbits = 6509
Recent Events:
The spacecraft continues to operate nominally in performing the beta supplement daily recording and transmission of science data. The mm028 sequence executed successfully from 00-97 (4/6/00) through 00-100 (4/9/00). The mm029 sequence also successfully executed from 00-101 (4/10/00) through 00-103 (4/12/00). The mm030 sequence was successfully uplinked on 00-102 (4/11/00) and will begin execution on 00-104 (4/13/00).
Included in the mm028 sequence, was the fifth and last star scan of the current MOC focus campaign, and a 24-hour period of MOLA off-nadir scans of the North and South Poles, on 00-99 through 00-100 (4/8-4/9/00). The spacecraft performance during the final MOC star scan was nominal, but the stars were not seen in the image. Thus two of the five scans resulted in not seeing the stars. All indications are that the missed stars were due to a knowledge error in designing the scans. A retry of the two missing star scans will be scheduled for later in April, with an updated scan attitude.
Spacecraft performance during the MOLA scans was also nominal, with a single STAREX (star processing code) anomaly occurring proceeding the first MOLA off-nadir orbit. A STAREX autonomous reset occurred shortly after the sequence commanded the IMU to high rate mode and issued an intentional bias reset prior to the first off-nadir slew. After the autonomous reset STAREX converged properly and performed nominally for the duration of the MOLA scans.
Spacecraft Health:
All subsystems are reporting nominal health.
There have been 16 uplinks to the spacecraft during the last week, including new star catalogs and ephemeris files, instrument command loads, and the mm030 and mm031 sequences. Total command files radiated to the spacecraft since launch is 4581.
Upcoming Events:
Planning for the replacement of the monthly fixed-HGA periods for the purpose of obtaining radio science atmospheric occultation egress data, prohibited in normal beta supplement operations due to HGA gimbal constraints, proceeds. The first Radio Science occultation egress scans are scheduled for 5/20-5/22/00. Work has been underway over the last month on determining the feasibility of performing a bistatic radar experiment, in which the Spacecraft signal will be reflected off the Martian surface to Earth. All design hurdles have been cleared and work is now underway to implement this experiment for a May 14 opportunity.